### v2

# Identifying optimal magnetic field configurations for decoherence mitigation of boron vacancies in hexagonal boron nitride

Abstract
The negatively charged boron vacancy center in 2D hexagonal boron nitride has emerged as a promising quantum sensor. However, its sensitivity is constrained due to ubiquitous nuclear spins in the environment. The nuclear spins, hyperfine coupled with the central electron spin, effectively behave as magnetic field fluctuators, leading to rapid decoherence. Here, we explore the effectiveness of static magnetic field strength and orientation in realizing peculiar subspaces that can lead to enhanced spin coherence. Specifically, using detailed numerical simulations of the spin Hamiltonian, we identify specific field configurations that minimize energy gradients and, consequently, are expected to facilitate decoherence suppression. We also develop an approximate analytical model based on the perturbation theory that accurately predicts these low-gradient subspaces for magnetic fields aligned with the electron spin quantization axis, applicable not only to boron vacancies but to any spin-1 electronic system coupled to nearby nuclear spins. Furthermore, to stimulate experimental validation, we estimate coherence lifetimes as a function of various bias field configurations and demonstrate that significant decoherence suppression can indeed be achieved in certain regions. These findings and the developed methodology offer valuable insights for mitigating decoherence in a low-field regime.
I Introduction
Optically active spin defects are gaining significant attention due to their potential in various quantum applications [1, 2, 3], including quantum sensing [4, 5, 6, 7, 3, 8], quantum computing [9, 10, 11, 12, 13], and quantum communication [14, 15]. A key feature of these defects is their extraordinary sensitivity to external perturbations such as magnetic fields [4, 5, 6, 7, 8, 16], electric fields [17, 18], strain [17, 19], and temperature [20, 21]. This sensitivity, coupled with the ability to perform optical readout [20, 22] of spin states at room temperature, enables the development of ultrasensitive nanoscale sensors. Among the most extensively studied systems is the nitrogen-vacancy ($NV$) center [23, 24, 25, 26] in diamond and spin defects in silicon carbide (SiC) [27, 28], which have paved the way for practical implementations in quantum sensing and metrology.
Despite their success, spin defects in three-dimensional (3D) materials such as diamond and SiC face challenges that limit their applicability. Specifically, their confinement within the bulk structure prevents spin probes from approaching the target sample with sub-nanometer resolution. In contrast, spin defects in two-dimensional (2D) materials offer a promising solution. Their atomic-scale thickness allows closer proximity to the sample [29, 30], enhancing spatial resolution. Moreover, 2D defects can be seamlessly integrated with photonic nanostructures [30, 31, 32, 33, 34] enabling the development of compact, on-chip quantum technologies [35, 36].
Among the spin defects in 2D materials [3], the negatively charged boron vacancy ($V_{B}^{-}$) in hexagonal boron nitride (hBN) [37, 38, 20, 39] has recently emerged as a leading contender. hBN, a Van der Waals material, provides a robust and tunable platform for hosting spin defects. The $V_{B}^{-}$ defect has demonstrated remarkable potential due to its optical and spin properties [22, 20], making it a promising candidate for high-resolution quantum sensing [40, 41, 19, 42]. Both theoretical and experimental [43, 44, 45] studies have highlighted the potential of $V_{B}^{-}$ in applications ranging from nanoscale magnetometry to temperature sensing.
$A_{xx}$ $A_{yy}$ $A_{zz}$ $A_{xy}$ $Q_{xx}$ $Q_{yy}$ $Q_{zz}$ $Q_{xy}$
${}^{14}N_{1}$ 46.944 90.025 48.158 0 -0.46 0.98 -0.52 0
${}^{14}N_{2}$ 79.406 58.170 48.159 -18.391 0.62 -0.1 -0.52 -0.623
${}^{14}N_{3}$ 79.406 58.170 48.159 18.391 0.62 -0.1 -0.52 0.623
Table 1: The table outlines all the hyperfine and quadrupolar tensor elements used in the numerical calculations.
A key goal within quantum sensing research is to enhance sensitivities through increasing the coherence time $(T_{2})$ of the probe system. In solid state systems the primary source of decoherence is magnetic field fluctuations in the spin bath surrounding the defect. On the materials engineering side, isotopic purification has been a successful strategy in systems like diamond [46] and SiC [47], however, hBN naturally lacks this avenue for isotopic refinement because of its chemical composition [48]. Therefore, alternative strategies to suppress decoherence are crucial to realize the full potential of $V_{B}^{-}$ defects in hBN.
Besides conventional dynamical decoupling methods [49, 50], a promising approach to address decoherence is identifying subspaces of the system’s Hilbert space that are minimally affected by spin fluctuations [51, 52, 53, 54, 55, 56, 57] in the bath. This requires a detailed understanding of how external parameters, such as magnetic field, influence the defect’s energy level structure and spin-transition properties.
In this study we investigate influence of the strength and the orientation of the static magnetic field on the transition energies of $V_{B}^{-}$ defect in hBN. We analyze energy gradient and curvature at low-bias magnetic fields ($<$ 25 mT), identifying regions where decoherence effects can be minimized. We show that these regions with minimal sensitivity to the magnetic fluctuation correspond to highly improved coherence and are critical for a variety of quantum sensing applications.
II System Hamiltonian
A missing boron atom in 2D-hBN creates a $V_{B}$ defect. In its negatively charged state, this defect hosts a spin-1 electronic system, surrounded by a spin-full nuclear bath, shown in schematic Fig. 1. The three $1^{st}$ shell ${}^{14}N$ nuclear spins (spin-1) are strongly coupled with the central $V_{B}^{-}$ electron spin (spin-1) via hyperfine interaction. The Hamiltonian of this composite system in the presence of an external magnetic field $\left(\bm{B}\right)$ can be written as:
{dmath}
H = D(S_z^2-23) + ϵ(S_y^2 - S_x^2) + γ_e B.S + ∑_i=1^3 S.A^i.I^i + ∑_i=1^3 γ_n^i B.I^i + ∑_i=1^3 I^i.Q^i.I^i. Here, $\bm{S}$, $\bm{I}$ are electron and nuclear spin operators, respectively. $D=3.45$ GHz is the zero field splitting (ZFS) term, $\epsilon$ is the strain term, which has been set to zero for this analysis, $\gamma_{e}$, $\gamma_{n}$ are the electron and nuclear gyromagnetic ratio, $\bm{A}$ is the hyperfine interaction tensor, $\bm{Q}$ is the quadrupole interaction tensor. The $z$-axis is the electron spin quantization axis, perpendicular to the hBN plane. Below, we outline the methodology and key steps involved in the numerical simulation of the composite $\left(V_{B}^{-},^{14}N\right)$ defect system.
FIG. 1: Structure of a $V_{B}^{-}$ defect in hBN, interacting with three nearest neighbor ${}^{14}N$ nuclear spin. Pink spheres represent boron atoms, blue spheres denote nitrogen atoms, and the red circle marks the boron vacancy site.
We utilize the quantum toolbox in Python (QuTiP) to construct spin operators for both electronic ($\bm{S}$) and nuclear ($\bm{I}$) subsystems. QuTiP’s built-in functions were employed to create the Hamiltonian matrix and compute the eigen-energies and eigen-states of the full spin system. The ability to handle high-dimensional Hilbert spaces and define tensor products of spin operators made QuTiP a natural choice for these calculations.
The hyperfine interaction tensor $\bm{A}$ for the first shell ${}^{14}N$ nuclear spins as outlined in Table 1, is obtained from the parameters reported by Gao et al. [58]. Since the first shell ${}^{14}N$ nuclear spins are arranged symmetrically around the central $V_{B}^{-}$ defect, the hyperfine interaction tensors of the other two nuclei can be obtained by rotating the tensor corresponding to the first nucleus about the symmetry axis ($y$-axis). The quadrupole interaction tensor $\bm{Q}$ for the first ${}^{14}N$ nuclear spin as given in Table 1, is based on the value reported by Gracheva et al. [59]. Using the similar rotational transformation used for hyperfine parameters, we derive the quadrupole tensors for the remaining two ${}^{14}N$ nuclei.
III Results and Discussion
FIG. 2: (a) Energy levels of the system as a function of the magnetic field applied parallel to the quantization axis of the electron spin. The black lines correspond to $m_{s}=0$, and orange and blue are for $m_{s}=\pm{1}$ electronic sub-level. These sub-levels split into 27 $(3\times 3\times 3)$ hyperfine levels. (b), (c) and (d) Energy levels as a function of the polar angle of the applied magnetic field with respect to the quantization axis for 1.7, 3.4, and 5.1 mT magnetic field, respectively. (e), (f) and (g) The energy gradient of a spin transition with the highest possible oscillator strength at the aforementioned fields.
Solid-state spin-based sensing protocols typically involve applying a bias magnetic field parallel to the electron spin dipole orientation. Here, we perform a comprehensive analysis of the electronic energy level structure of the composite $\left(V_{B}^{-},^{14}N\right)$ defect as a function of varied bias field amplitude and direction given by:
[…]
Here, $\theta$ and $\phi$ are polar and azimuthal angles, respectively, while $B_{0}$ is the field strength.
Firstly, Fig. 2(a) depicts the energy levels of the system for the static field that is applied along the $z$-axis. Here, the primary splitting observed between the $m_{s}=0$ and $m_{s}=\pm{1}$ states is due to the zero field interaction term. A smaller splitting between the $m_{s}=+1$ and $m_{s}=-1$ sublevels arises from the hyperfine interaction between the electron spin and the three $1^{st}$ shell ${}^{14}N$ nuclear spins. If the magnetic field is applied along the spin quantization axis, these nuclear spins can be considered equivalent, effectively giving rise to a spin-three system. As a result, the energy sublevels can be grouped into seven distinct hyperfine levels, as seen in Fig. 2(a). Additionally, the $m_{s}=+1$ and $m_{s}=-1$ levels demonstrate avoided crossings at relatively low fields due to the intricate interplay between Zeeman and hyperfine terms in the system’s Hamiltonian; the detailed discussion of this phenomenon is provided in the later part of this section.
Secondly, we investigate the energy level fluctuations as a function of magnetic field orientation at three distinct anti-crossing points. Figs. 2(b), 2(c), and 2(d) display the polar angle-dependent energy variations for field strengths 1.7, 3.4, and 5.1 mT. In this regime, the energy levels show minimal dependence on the magnetic field orientation, indicating that they are largely unaffected by local fluctuations.
To explore the behavior in detail, we vary both the polar angle $\left(\theta\right)$ and azimuthal angles $\left(\phi\right)$ while keeping the magnitude of the bias field $\left(B_{0}\right)$ constant. The workflow of gradient calculation is as follows:
  1. 1.
    At each point on the shell, defined by $\left(\theta,\phi\right)$ we introduce small fluctuations $\bm{b}$ in magnetic field $\bm{B}$. Hence, total field strength $\bm{B_{tot}}=\bm{B}\pm\bm{b}$.
  2. 2.
    We then find the eigen-energies of the full Hamiltonian and transition energy $\left(f\right)$ for all possible transitions, at each of these points.
  3. 3.
    Subsequently, we calculate gradient of the transition energy $\left(\partial f/\partial b_{x},\hskip 2.84544pt\partial f/\partial b_{y},\hskip 2.84544pt\partial f/\partial b_{z}\right)$ along each direction and finally, evaluate the magnitude of the gradient as follows:
    $|\nabla f|=\sqrt{\left(\frac{\partial f}{\partial b_{x}}\right)^{2}+\left(\frac{\partial f}{\partial b_{y}}\right)^{2}+\left(\frac{\partial f}{\partial b_{z}}\right)^{2}}.$ (2)
The gradient $|\nabla f|$ at each $\left(\theta,\phi\right)$ is mapped to a color scale to visualize the magnetic sensitivity of the transition energy across the sphere. Figs. 2(e), 2(f), and 2(g) illustrate the gradient of the transition energy for the most probable transitions with a transition probability of more than 30%. The spherical shell’s radius $\left(B_{0}\right)$ represents the magnetic field strengths at the level anti-crossing points, while the color gradient from blue to yellow signifies regions of low to high energy gradient, reflecting variations in magnetic sensitivity.
Our analysis reveals that the gradient at low field strengths is minimal at the poles. As the field strength increases, the gradient gradually shifts towards the equator, forming distinctive rings highly dependent on the bias field. This suggests that beyond this regime of a very low static field, the low gradient subspaces can be accessed via the application of the field in a transversal direction.
FIG. 3: (a) Energy levels as a function of the transverse magnetic field. The black lines correspond to $m_{s}=0$, and orange and blue are for $m_{s}=\pm{1}$ electronic sub-level. (b), (c) and (d) Energy levels as a function of the polar angle of applied magnetic field with respect to the quantization axis for 7.5, 15, and 23.5 mT magnetic field, respectively. (e), (f) and (g) The energy gradient of a spin transition with the highest possible oscillator strength at the aforementioned fields. The gradients are minimal near the equatorial region for the transverse fields.
In Fig. 3(a) we present the ground state energy levels for transverse applied magnetic fields. In this configuration, the primary splitting between $m_{s}=0$ and $m_{s}=\pm 1$ states remains dominated by crystal field interactions $(D)$. However, applying the transverse field leads to a significant mixing of the $m_{s}=+1$ and $m_{s}=-1$ spin states, significantly modifying the energy level landscape.
We conduct a magnetic field orientation-dependent analysis following the aforementioned approach. In this case, we investigate the energy levels at three distinct magnetic field strengths, namely 7.5, 15, and 23.5 mT, as shown in Figs. 3(b), 3(c), and 3(d), respectively. In this orientation, the $m_{s}=+1$ and $m_{s}=-1$ levels undergo anti-crossing, leading to the low gradient $|\nabla f|$, effectively making the spin states less susceptible to fluctuations in the applied field.
Figs. 3(e), 3(f), and 3(g) depict the energy gradient of a spin transition having highest oscillator strength within this field regime. As anticipated, the lowest gradient values occur near the transverse plane at higher fields. This finding suggests that applying a transverse magnetic field becomes increasingly effective in shielding the spin states from magnetic fluctuations as the field strength is increased, providing enhanced stability and protection for the spin system.
Fig. 4(a) shows the overall gradient of all possible transitions as a function of the applied magnetic field strength oriented parallel to the crystal field axis. At low parallel magnetic fields, the system exhibits notably low gradients at specific field strengths. The sharp dips in the gradient are essentially due to the competing interactions in the Hamiltonian, namely, Zeeman and electron-nuclear. To obtain an intuitive description of these features, we develop an approximate analytical approach based on the perturbation theory. Under a parallel bias magnetic field, the three first-shell ${}^{14}N$ nuclear spins interact identically to the $V_{B}^{-}$ spin. Consequently, the system can be effectively described as an interacting spin-1, spin-3 system. Under secular approximation, the Hamiltonian can be written as:
$H=DS_{z}^{2}+\gamma_{e}B_{z}S_{z}+A_{zz}S_{z}I_{z},$ (3)
where $B_{z}$ is bias magnetic field, D and $A_{zz}$ are parallel ZFS and hyperfine constant. At low bias field the perturbative contributions are considered to arise from three sources: the Zeeman interaction with bias field, Zeeman interaction with external magnetic fluctuation $\bm{b}$ and hyperfine interaction with first shell nuclear spins. Applying second order perturbation to the Hamiltonian, we obtain the total energy as:
$E_{|{m_{s},m_{I}}\rangle}=E_{|{m_{s},m_{I}}\rangle}^{0}+E_{|{m_{s},m_{I}}\rangle}^{1}+E_{|{m_{s},m_{I}}\rangle}^{2}.$ (4)
The unperturbed energy for the states $|{m_{s},m_{I}}\rangle$,
[…]
The first-order energy correction is given by,
[…]
The second-order energy correction is expressed as,
[…]
FIG. 4: Energy gradient and curvature associated with all $(2\times 27\times 27)$ transitions between $m_{s}=0$ and $m_{s}=\pm{1}$ as a function of the magnetic field applied parallel (a) and (c), perpendicular(b) and (d) to the quantization axis. The color bar corresponds to the magnitude of the gradient or the curvature; values beyond an arbitrary threshold are whitened for visibility purposes.
Thus, the approximate transition energy of transitions satisfying condition, $\Delta m_{s}=\pm 1,\Delta m_{I}=0$ is given by (in-depth calculations are provided in the Appendix VII):
$f=D\pm\sqrt{\left[\gamma_{e}\left(B_{z}+b_{z}\right)+m_{I}A_{zz}\right]^{2}\pm\left[\gamma_{e}\left(b_{x}^{2}+b_{y}^{2}\right)\right]^{2}}.$ (6)
The amplitude of the gradient of these transition energies can be obtained using Eq.(2), where, the partial derivatives of the transition frequency with respect to $b_{x}$, $b_{y}$ , and $b_{z}$ are as follows:
$\begin{split}\frac{\partial f}{\partial b_{x}}&=\frac{2m_{s}^{2}\gamma_{e}^{2}b_{x}}{Y},\quad\frac{\partial f}{\partial b_{y}}=\frac{2m_{s}^{2}\gamma_{e}^{2}b_{y}}{Y},\\ \frac{\partial f}{\partial b_{z}}&=\frac{m_{s}\gamma_{e}\left[\gamma_{e}\left(B_{z}+b_{z}\right)+m_{I}A_{zz}\right]}{Y}.\end{split}$ (7)
here, $Y=\sqrt{\left[\gamma_{e}\left(B_{z}+b_{z}\right)+m_{I}A_{zz}\right]^{2}+m_{s}\left[\gamma_{e}\left(b_{x}^{2}+b_{y}^{2}\right)\right]^{2}}.$ Assuming, $\gamma_{e}B_{z},A_{zz}>>\gamma_{e}b_{x},\gamma_{e}b_{y},\gamma_{e}b_{z}$, one can find the condition for minimal gradient as follows:
$|{\nabla{f}}|=0\implies\gamma_{e}B_{z}+m_{I}A_{zz}=0,$ (8)
$\boxed{B_{z}=-\frac{m_{I}A_{zz}}{\gamma_{e}}}$ (9)
For the systems comprising multiple nuclear spin shells, Eq. (9) will get extended to $B_{z}=-\sum_{i}\frac{m_{I_{i}}A_{zz_{i}}}{\gamma_{e}},$ where $i$ represents the nuclear spin shell index. If the defect is experiencing out-of-plane interactions due to the presence of multiple hBN layers, the hyperfine coupling $A_{zz}$ needs to be replaced with $\sqrt{A{zx}^{2}+A{zy}^{2}+A{zz}^{2}}$. The matching condition presented in Eq. 9 will have to be amended to account for the $\theta$ dependence of the Zeeman term, if the applied field is not aligned with the quantization axis. However, it is apparent that, as the magnetic field strength increases, the matching between Zeeman and hyperfine splittings occurs at larger $\theta$ values. Consequently, the avoided crossings will shift to larger $\theta$ values for higher magnetic field magnitudes as depicted in Figs. 2 and 3. Similar phenomena have been observed in nitrogen vacancy centers in diamond [57].
Furthermore, we analyze the specific scenario when the magnetic field is applied perpendicular to the crystal field axis. Fig. 4(b) shows that the transition gradients are minimal at zero field, however, their rate of increase as a function of the field strength is dramatically suppressed for the transversal field. A lower gradient indicates a reduced first-order response to external perturbations, implying that the system is less affected by small magnetic fluctuations. In such cases, the second-order effect, characterized by the curvature of the transition energy, becomes more relevant. Therefore, we compute the mean curvature for parallel and transverse field variations. The procedure to compute the curvature is as follows:
  1. 1.
    At each magnetic field, we introduce small fluctuations in $18$ possible directions: $6$ along the principal axes and $12$ pointing towards the diagonal of the $xy$, $xz$, and $yz$ planes.
  2. 2.
    We then calculate the eigen-energies of the full Hamiltonian and transition energy $(f)$ for all possible transitions.
  3. 3.
    Subsequently, we calculate first order $(f_{x},f_{y},f_{z})$, second order $(f_{xx},f_{yy},f_{zz})$ and mixed $(f_{xy},f_{xz},f_{yz)}$ derivatives of the transition energy at each magnetic field and finally, evaluate the magnitude of mean curvature as follows [60]:
    $\begin{split}K_{M}&=\frac{\nabla f\times H(f)\times\nabla f^{T}-\left|\nabla f\right|^{2}Tr(H)}{2\left|\nabla f\right|^{3}}\\ &=\frac{1}{2\left|\nabla f\right|^{3}}|f_{xx}\left(f_{y}^{2}+f_{z}^{2}\right)+f_{yy}\left(f_{x}^{2}+f_{z}^{2}\right)+f_{zz}\\ &\left(f_{x}^{2}+f_{y}^{2}\right)-2\left(f_{xy}f_{x}f_{y}+f_{xz}f_{x}f_{z}+f_{yz}f_{y}f_{z}\right)|\end{split}$ (10)
    Here, $H(f)$ is the Hessian matrix corresponding to transition energy $(f)$.
FIG. 5: Estimated coherence lifetime as a function of bias magnetic field for (a) parallel field configuration, and (b) transverse field configuration. Here, the upper bound of the colorbar has been restricted to 1 $\mu$s for enhancing the visibility of the high $T_{2}$ regions. Dependence of $T_{2}$ as a function of the applied field corresponding to select transition energies for (c) parallel field configuration, and (d) transverse field configuration. The red arrows in (a) and (b) indicate the chosen energies, and the gray lines in (c) indicate the magnetic field values that satisfy Eq. 9.
The curvature (i.e., second order effect due to magnetic fluctuations) of transition energies for fields applied in parallel and transverse orientations are shown in Figs. 4 (c) and 4(d) respectively. As anticipated from the geometry of the problem, the curvature is larger at the anti-crossing points compared to the other field regions. Besides the anti-crossing points, the curvature increases as a function of the field strength for the parallel field, nonetheless, the magnitude and variations are considerably lower as compared to the transverse field scenario. Moreover, for the transversal field application, although the magnitude of the curvature is frequency dependent, it decreases with magnetic field strength.
Although our analysis does not explicitly facilitate evaluation of the exact $T_{2}$ values, we implement the following route to illustrate coherence enhancement corresponding to the investigated scenarios. Assuming that $T_{2}$ is inversely proportional to the frequency broadening caused by the magnetic field perturbations, its dependence on energy gradients and curvatures can be expressed as:
$T_{2}\approx\frac{1}{\sqrt{\left(\frac{df}{dB}\sigma_{B}\right)^{2}+\frac{1}{2}\left(\frac{d^{2}f}{dB^{2}}\sigma_{B}^{2}\right)^{2}}}$ (11)
where $\sigma_{B}$ is the standard deviation of the external field noise (assuming the normal distribution). Here, we assume that magnetic noise is the leading source of decoherence, which is the case for $V_{B}^{-}$ in hBN due to the nuclear spin dominated bath. Taking into consideration the reported values $T_{2}$, both experimental as well as those obtained via Cluster Correlation Expansion (CCE) [61], we estimate $\sigma_{B}$ to be 388.5 $\mu T$ at 15 $mT$. This allows for the calculation of $T_{2}$ associated with each of the transition energies at various field configurations using Eq. 11.
The color plots of $T_{2}$ for the parallel and the perpendicular magnetic field scenarios are presented in Figs. 5(a) and (b). Moreover, the cross-sections taken at specific transition energies are shown in Figs. 5(c) and (d). Here, the cross-sections have a width of 10 $MHz$ on each side of the central transition energy. For the parallel fields, we obtain extremely long coherence lifetimes at the peculiar points that satisfy Eq. 9. Additionally, some extra features are observed at frequencies where the approximate analytical solution becomes inadequate to capture the complex energy level structure. The linewidths of the individual $T_{2}$ peaks are of the order of a few tens of $\mu T$, thus requiring a relatively high control over the bias field for the practical realization. On the other hand, the transverse field configurations appear to be much less sensitive to the bias field amplitude, though the $T_{2}$ enhancement factor is not as high as the one observed for the parallel cases. In an actual experiment, we expect the enhanced $T_{2}$ features to be less pronounced due to the broadening effects arising from the factors such as interactions with the rest of the bath, bias field inhomogeneities, and the finite bandwidth of the experimental probe pulse. The oscillator strengths of the probed transitions will also influence the outcome of the experiment. Nonetheless, we anticipate noticeably higher coherence lifetimes close to the predicted configurations.
IV Summary & Outlook
Our comprehensive investigation of magnetic field dependent properties of the composite $\left(V_{B}^{-},^{14}N\right)$ defect uncovers rich physics. In particular, by performing an in-depth numerical simulation of the system’s spin Hamiltonian, we identify distinct anti-crossing points, corresponding to specific field strengths and orientations where the energy gradients reach their minimum value. Furthermore, we derive an approximate analytical expression to estimate conditions under which gradient minimization occurs when the magnetic field is oriented along the quantization axis. This model is widely applicable as the low-gradient regions can be evaluated for an arbitrary spin-1 electronic system coupled to an arbitrary nuclear spin. Additionally, we compute curvatures of the transition energies that can play an important role in certain low-gradient regions. Finally, we estimate experimentally observable $T_{2}$ values and demonstrate that coherence can be improved substantially if specific field configurations are realized. More generally, the analysis presented in this study will serve as a guide for interpreting outcomes of magnetic field dependent $T_{2}$ measurements for a variety of systems and will stimulate an investigation of such coherence preserving configurations using more precise numerical techniques as CCE.
Our results have significant implications for low-field quantum sensing, where decoherence mitigation remains a critical challenge. Specifically, the identified gradient dips and field-insensitive regions highlight potential pathways for improving spin coherence in $V_{B}^{-}$ defects, thus advancing the development of quantum sensing applications such as electometry and thermometry, based on 2D materials.
V Acknowledgments
We are grateful to Professor John J. L. Morton for his valuable insights. S.D. thanks the Indian Institute of Technology, Madras, India, and the Science and Engineering Research Board (SERB Grant No. SRG/2023/000322), India, for start-up funding. S.D. and B.M. acknowledge the use of facilities supported by a grant from the Mphasis F1 Foundation given to the Center for Quantum Information, Communication, and Computing (CQuICC).
VI Code Availability
The code for Hamiltonian diagonalization as well as gradient and curvature evaluation is available here.
VII Appendix: Secular Hamiltonian and Perturbation Analysis
The secular Hamiltonian for the system under consideration is given by
$H=DS_{z}^{2}+\gamma_{e}B_{z}S_{z}+A_{zz}S_{z}I_{z},$ (A1)
where $B_{z}$ is the bias magnetic field, $D$ is the zero-field splitting (ZFS) term, and $A_{zz}$ is the hyperfine coupling constant. The unperturbed Hamiltonian, $H^{0}$ is defined as
$H^{0}=DS_{z}^{2}.$ (A2)
This term represents the dominant energy contribution due to the ZFS, which splits the $m_{s}=0$ and $m_{s}=\pm 1$ spin states of the system. The perturbation Hamiltonian, $H^{1}$, accounts for contributions from magnetic field and hyperfine interaction, is expressed as {dmath} H1= γe[bxSx+ bySy+ (Bz+ bz)Sz]+ AzzSzIz= γe[bx(S++ S-2)+ by(S+- S-2i)+(Bz+ bz)Sz]+ AzzSzIz. Here, $b_{x},b_{y},b_{z}$ are small fluctuations in the magnetic field components, $S_{+}=S_{x}+iS_{y}$ and $S_{-}=S_{x}-iS_{y}$ are the spin-raising and lowering operators. The unperturbed energy levels corresponding to $H^{0}$ are given by
$E_{|{m_{s},m_{I}}\rangle}^{0}=Dm_{s}^{2},$ (A3)
The states $|{m_{s}=0,m_{I}}\rangle$ are degenerate at zero energy, whereas the states $|{m_{s}=\pm 1,m_{I}}\rangle$ are degenerate at energy $D$. To account for the effects of $H^{1}$, first-order perturbation theory is applied. The first-order corrections to the Hamiltonian for $m_{s}=\pm 1$ are given by the perturbation matrix $W^{\pm 1}=[w]_{14\times 14}$, where
[…]
$w^{0}_{m_{I},m_{I}}$ $=\braket{0,m_{I}}{H^{1}}{0,m_{I}}=0,$ (A5a)
$w^{0}_{m_{I},m_{I}^{^{\prime}}}$ $=\braket{1,m_{I}}{H^{1}}{1,m_{I}^{^{\prime}}}=0.$ (A5b)
This indicates that there are no first-order energy corrections for states with $m_{s}=0$. The first-order perturbation energy is given by:
$E_{|{m_{s},m_{I}}\rangle}^{1}=m_{s}\gamma_{e}B_{z}+m_{s}m_{I}A_{zz}.$ (A6)
The first-order perturbation lifts the degeneracy of $|{m_{s}=\pm 1,m_{I}}\rangle$ states, but the $|{m_{s}=0,m_{I}}\rangle$ states remain degenerate. For the second-order perturbation, the energy correction is expressed as:
$E_{|{m_{s}=\pm 1,m_{I}}\rangle}^{2}$ $=\sum_{n\neq m}\frac{|{\braket{\psi_{m}}{H^{1}}{\psi_{n}}}|^{2}}{E_{n}-E_{m}}$ (A7)
=±$=\frac{\pm\gamma_{e}^{2}\left(b_{x}^{2}+b_{y}^{2}\right)}{2\left[\gamma_{e}\left(B_{z}+b_{z}\right)+m_{I}A_{zz}\right]}.$ (A8)
The second order perturbation matrix for degenerate states corresponding to $m_{s}=0$ is given as:
$\Lambda^{0}=\left[\lambda\right]_{7\times 7},$ (A9)
where
$\lambda_{ij}^{0}=\sum_{n\neq m}\frac{\braket{\phi_{n,i}}{H^{1}}{\psi_{m}}\braket{\psi_{m}}{H^{1}}{\phi_{n,j}}}{E_{n}^{(0)}-E_{m}^{(0)}}=0.$ (A10)
Here, $\phi_{n}$ are degenerate states, and $\psi_{m}$ are all other states. The second-order energy correction corresponding to $m_{s}=0$ is $E_{|{m_{s}=0,m_{I}}\rangle}^{2}=0$. Now, the general expression for second-order perturbation energy is expressed as:
$E_{|{m_{s},m_{I}}\rangle}^{2}=\frac{m_{s}^{2}\gamma_{e}^{2}\left(b_{x}^{2}+b_{y}^{2}\right)}{2\left[m_{s}\gamma_{e}\left(B_{z}+b_{z}\right)+m_{s}m_{I}A_{zz}\right]}.$ (A11)
The total energy of the system reads:
$E$ $=E^{0}+E^{1}+E^{2}$ (A12)
$=m_{s}^{2}D+m_{s}Y,$
where,
$Y=\sqrt{\left[\gamma_{e}\left(B_{z}+b_{z}\right)+m_{I}A_{zz}\right]^{2}+m_{s}\left[\gamma_{e}\left(b_{x}^{2}+b_{y}^{2}\right)\right]^{2}}.$ (A13)
The transition frequency is defined under the condition $\Delta m_{s}=\pm 1$ and $\Delta m_{I}=0$, and is given by:
$f=D\pm\sqrt{\left[\gamma_{e}\left(B_{z}+b_{z}\right)+m_{I}A_{zz}\right]^{2}\pm\left[\gamma_{e}\left(b_{x}^{2}+b_{y}^{2}\right)\right]^{2}}.$ (A14)
The partial derivatives of the transition frequency with respect to $b_{x}$, $b_{y}$, and $b_{z}$ are as follows:
$\frac{\partial f}{\partial b_{x}}=\frac{2m_{s}^{2}\gamma_{e}^{2}b_{x}}{\sqrt{\left[\gamma_{e}\left(B_{z}+b_{z}\right)+m_{I}A_{zz}\right]^{2}+m_{s}\left[\gamma_{e}\left(b_{x}^{2}+b_{y}^{2}\right)\right]^{2}}},$ (A15)
$\frac{\partial f}{\partial b_{y}}=\frac{2m_{s}^{2}\gamma_{e}^{2}b_{y}}{\sqrt{\left[\gamma_{e}\left(B_{z}+b_{z}\right)+m_{I}A_{zz}\right]^{2}+m_{s}\left[\gamma_{e}\left(b_{x}^{2}+b_{y}^{2}\right)\right]^{2}}},$ (A16)
and
$\frac{\partial f}{\partial b_{z}}=\frac{m_{s}\gamma_{e}\left[\gamma_{e}\left(B_{z}+b_{z}\right)+m_{I}A_{zz}\right]}{\sqrt{\left[\gamma_{e}\left(B_{z}+b_{z}\right)+m_{I}A_{zz}\right]^{2}+m_{s}\left[\gamma_{e}\left(b_{x}^{2}+b_{y}^{2}\right)\right]^{2}}}.$ (A17)
The gradient of the transition energies is then given by:
$\frac{df}{db}=\sqrt{\left(\frac{\partial f}{\partial b_{x}}\right)^{2}+\left(\frac{\partial f}{\partial b_{y}}\right)^{2}+\left(\frac{\partial f}{\partial b_{z}}\right)^{2}}.$ (A18)
The gradient is minimized when:
$\frac{df}{db}=0,\implies\gamma_{e}B_{z}+m_{I}A_{zz}=0.$ (A19)
Under the condition of small perturbation $i.e.,$ $\gamma_{e}B_{z},A_{zz}>>\gamma_{e}b_{x},\gamma_{e}b_{y},\gamma_{e}b_{z}$, the bias magnetic field $B_{z}$ satisfies:
$B_{z}=-\frac{m_{I}A_{zz}}{\gamma_{e}}.$ (A20)